OpenStudy (anonymous):

How do I know which is right- y2-y1 over x2-x1 or y1-y2 over x1-x2 People say it's the same on yahoo answers but they get different answers a lot. HELP It's different. Let's say (-2,8) and (2,-1). First way: -1 - 8 = -9 2 - (-2) = 0 So... -9/0 Second way: 8- (-1) = 9 -2 - 2 = 0 So... 9/0 See it's like this all the time. In this example they are the same because it is undefined, but in other examples they are regular numbers that are negatives.

OpenStudy (anonymous):

Its the same, but in different order.

OpenStudy (anonymous):

either one works... so long as your going the same way. point 1 to point 2 or point 2 to point 1

OpenStudy (anonymous):

check it out... \[ \frac{y_{2} - y_{1}}{x_{2} - x_{1}}=\frac{-(y_{1} - y_{2})}{-(x_{1} - x_{2})}=\frac{y_{1} - y_{2}}{x_{1} - x_{2}}\]

OpenStudy (psymon):

If you choose a y1, the x1 better be from the same point. As long as you make x1,y1 the same point and x2,y2 the same point, it really doesnt matter Ex. (3, -4) and (-5,2) So lets make the (-5,2) my point 2 \[\frac{ 2-(-4) }{ -5-3 }=\frac{ -3 }{ 4 } \] Now lets do it the other way, Ill have (3,-4) be my point two: \[\frac{ -4-2 }{ 3-(-5) }=\frac{ -3 }{ 4 }\]

OpenStudy (anonymous):

yes... check your arithmetic @Mia_Rachel

OpenStudy (anonymous):

What did I do wrong?

OpenStudy (ness9630):

I think you guys scared her

OpenStudy (anonymous):

2-(-2) = 2 + 2 because subtraction is addition of the opposite

OpenStudy (ness9630):

Psymon and pg scaring users with their math? Not surprising

OpenStudy (anonymous):

also, -2 - 2 = -2 +(-2) = -4

OpenStudy (anonymous):

what? when did math get scary?

OpenStudy (psymon):

\[\frac{ -1-8 }{ 2-(-2) }=\frac{ -9 }{ 4 } \] \[\frac{ 8-(-1) }{ -2-2 }=\frac{ -9 }{ 4 } \]

OpenStudy (anonymous):

challenging, frustrating, down right nasty... but not scary!

OpenStudy (psymon):

I think for a lot of people it gets scary when there are no numbers. Which is kind of ironic.

OpenStudy (anonymous):

letters are so much easier to deal with...

OpenStudy (anonymous):

I'm sorry it's not (2,-1) it's (-2,-1).

OpenStudy (anonymous):

in that case you're dividing by 0, either way you do it. since division by 0 leads to an undefined quotient, the slope is undefined in both cases. you'll learn more about this in calculus when you actually look closer at division by 0.

OpenStudy (psymon):

Well, either way, even though the numbers I used were not the correct numbers, can you see how you still get the same answer if you do the math correct?

OpenStudy (anonymous):

Can you do it with the same numbers too? I like to see examples. :/

OpenStudy (psymon):

I gave an example way up above already using numbers I invented.

OpenStudy (anonymous):

With the numbers I use please. (-2,8) (-2,-1)

OpenStudy (anonymous):

what you did is correct... you get 2 seemingly different answers, but they're not. in both cases you're dividing by 0 which is undefined... meaning both ways actually give you an undefined slope. if you examine the numbers, they have the same x coordinate meaning you have a vertical line. what's the slope of a vertical line?

OpenStudy (anonymous):

undefined

OpenStudy (psymon):

Honestly, your example doesnt perfectly show how the signs end up being the same. If one of those -2's were any different number, you would come out with the same answer and you would not have to worry about having the incorrect sign. In your case it's just undefined no matter what, which means a vertical line.

OpenStudy (anonymous):

Thanks! Sorry, I just got back to school and I can't remember anything in my math class (Alg II/Trig) and I'm in 9th grade.

OpenStudy (anonymous):

But can someone solve it for me? Apparently the answer is x=-2 (back of textbook) but can someone show me step by step how that happens?

OpenStudy (anonymous):

exactly... and what is \[\frac{9}{0} \text{ ?}\] what about \[-\frac{9}{0} \text{ ?}\]

OpenStudy (anonymous):

they're both undefined, right? like i said, you'll look more closely at this in calculus.

OpenStudy (anonymous):

Someone please show me step by step :(

OpenStudy (anonymous):

for the whole problem

OpenStudy (anonymous):

you already did it Mia... what is giving you trouble?

OpenStudy (anonymous):

I mean make up the linear equation with that. I still need to know b (y=mx+b).

OpenStudy (anonymous):

check it out... you have a vertical line. they have a special form (as does a horzontal line). for a vertical line, it's x = some number. in your case, x = -2 because no matter what y is, x is always -2. for a horizontal line, the equation will always be y = some number, for example, y = 5. no matter what x is, y will always be 5 

OpenStudy (anonymous):

I know that, but where did you get the -2?

OpenStudy (anonymous):

from the points you gave... (-2, 8) and (-2, -1) see, x is always -2.

OpenStudy (anonymous):

OH DUHHHHHHH

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

But what's the mathematical way you find that?

OpenStudy (anonymous):

when you find that the slope is undefined, that's when you know you have a vertical line. then you know the eqution of the line willl be in that form. likewise, when you get a slope of 0, you know you have a horizontal line and that the equation will be y = some number and some number will be the y coordinate of any of your points.

OpenStudy (anonymous):

You are a SAVIOR. THANKS I'm done with all of my questions xD

OpenStudy (anonymous):

thanks for sticking with it and making sure you understand! you're on your way to being an excellent mathematician!!!

OpenStudy (anonymous):

Ahaha thanks, I'm stubborn (:

OpenStudy (anonymous):

an excellent quality for anyone who wants to learn and understand ; )